OpenStudy (anonymous):

identify a pattern in each arithmetic seqences find the missing terms 12 24 36 48 ( ) ( ) ()

jimthompson5910 (jim_thompson5910):

We're adding 12 to each term. Notice how 36-24 = 12 and 48 - 36 = 12 So the term after 48 is 48+12 = 60 And the term after 60 is 60+12 = 72 What's the next term?

OpenStudy (anonymous):

84

jimthompson5910 (jim_thompson5910):

Good, you got it.

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

can u help me n a nother

jimthompson5910 (jim_thompson5910):

sure